# Data Fusion for Heterogeneous Treatment Effect Estimation with Multi-Task Gaussian Processes

### v1

@@ Abstract
Bridging the gap between internal and external validity is crucial for heterogeneous treatment effect estimation. Randomised controlled trials (RCTs), favoured for their internal validity due to randomisation, often encounter challenges in generalising findings due to strict eligibility criteria. Observational studies on the other hand, provide external validity advantages through larger and more representative samples but suffer from compromised internal validity due to unmeasured confounding. Motivated by these complementary characteristics, we propose a novel Bayesian nonparametric approach leveraging multi-task Gaussian processes to integrate data from both RCTs and observational studies. In particular, we introduce a parameter which controls the degree of borrowing between the datasets and prevents the observational dataset from dominating the estimation. The value of the parameter can be either user-set or chosen through a data-adaptive procedure. Our approach outperforms other methods in point predictions across the covariate support of the observational study, and furthermore provides a calibrated measure of uncertainty for the estimated treatment effects, which is crucial when extrapolating. We demonstrate the robust performance of our approach in diverse scenarios through multiple simulation studies and a real-world education randomised trial.
@@ 1 Introduction
Treatment effect estimation is an important task in many applications, including medicine, epidemiology, and the social sciences. The goal is to quantify the expected impact of a particular intervention in a given target population. For example, in medicine, one may seek to estimate the expected benefit or harm from taking a particular drug relative to placebo for a patient with a certain disease.
The Average Treatment Effect (ATE) from a study is often used as evidence for whether a particular intervention should be applied to an individual or unit in a given target population. The reliability of this evidence rests on two key assumptions. Firstly, the study population is assumed to be sufficiently similar to the target population, so that an estimate of the treatment effect is applicable to the target population in which the intervention under investigation will eventually be applied. Secondly, the average treatment effect is assumed to be close to an individual’s treatment effect.
This second assumption of homogeneity across individuals is often violated: the effect of an intervention within a population may vary systematically with respect to a particular covariate or combination of covariates. Returning to our medical example, a drug may have different effects for an individual as a function of their age or sex. In these cases, understanding the heterogeneity of treatment effects is critical to ensure that appropriate decisions are taken for all individuals, not just those who are similar to the average or belong to pre-specified subgroups that poorly reflect heterogeneity within the population (Brantner et al., 2023).
Whether the results of a study are relied upon for a particular decision or policy also depends on the statistical significance of the corresponding estimates. A proposed intervention with an estimated positive treatment effect may not be adopted if statistical variability indicates that this estimate is not significantly away from zero. Thus, an appropriate characterisation of the uncertainty of an estimate is also required. While this is reasonably straightforward in the computation of average treatment effects, uncertainty quantification for heterogeneous treatment effects is more challenging, especially when making inferences about individuals not represented in the randomised controlled trial (RCT), where classical methods rely heavily on extrapolation without proper variance inflation to reflect uncertainty (Degtiar and Rose, 2023).
The gold standard for treatment effect estimation remains the RCT. The random allocation of treatment to individuals in the study guards against the effect of confounding, assuming perfect compliance with allocated treatment. As a result, under a few additional assumptions, the difference in average outcomes between individuals in the treatment and control arms is unbiased for the average treatment effect in the study population.
There are, however, drawbacks to the use of RCTs (Frieden, 2017). They are typically expensive to carry out, and sample sizes can be limited as a result. While they may be powered enough to detect the main effect, they are typically underpowered for subgroup effects. Further, the study population may not be representative of the target population. This can be for a variety of reasons such as inclusion/exclusion criteria of individuals into the study (either on statistical, practical, or ethical grounds), or selection bias of individuals into the study; for instance, there may be differential acceptance to “invitation to participate" from eligible candidates according to, for example, socioeconomic status, educational levels or ethnicity (National Academies of Sciences, Engineering, and Medicine, 2022).
@@
Observational, or ‘real-world’ studies, provide an alternative source of data for treatment effect estimation. In contrast to RCTs, they generally benefit from larger sample sizes and tend to be more representative of the target population. However, a number of assumptions must be made to ensure the validity of treatment effect estimations from observational data. In particular, one must assume that there are no unmeasured confounders - hidden variables that affect both the outcome of interest and the treatment allocation. While it may be possible to adjust for the effect of observed confounders, there is no way to mitigate the effect of unmeasured confounding without making strong assumptions about the nature of confounding.
@@
The complementary nature of RCT and observational data indicates the potential benefits from combining the two sources to obtain better heterogeneous treatment effect estimates. Such data fusion approaches have received increased interest in recent years, and a range of methods have been developed to integrate information from multiple sources of data, including a limited number of applications to causal inference problems (see Colnet et al. (2023) for a recent review). These typically rely on strong, untestable assumptions about the structure of the confounding effect and, moreover, are not able to provide uncertainty quantification around the point estimates of heterogeneous treatment effects.
In this paper, we introduce a Bayesian non-parametric approach to obtain estimates of heterogeneous treatment effects for the target population, with uncertainty quantification. Our approach is based on multi-output Gaussian Processes, where the limited RCT observations are used to correct the confounding effect on the observational data. Our key contributions are as follows:
  • We introduce an Intrinsic Coregionalization Model (ICM) for heteregeneous treatment effect estimation, which we call Causal-ICM. Causal-ICM is a multi-task Gaussian process (GP) model that enables accurate treatment effect estimation within and outside of the support of the RCT and calibrated uncertainty quantification of treatment effect estimates across the support of the target population, without relying on strong assumptions on the structure of the hidden confounding effect.
  • Causal-ICM has a straightforward interpretation, relying on a key hyperparameter that quantifies the measure of trust in the quality of the observational data, and can be tuned data-adaptively. We provide theory to demonstrate that the ICM is robust against the dominating nature of a large observational dataset.
  • We demonstrate excellent empirical performance in simulations and a real dataset relative to other benchmarks for both point estimation and uncertainty quantification.
@@ 2 Related Literature
The challenge of heterogeneous treatment effect (HTE) estimation has received renewed interest in recent years. This is typically approached via the conditional average treatment effect (CATE) function, which quantifies the expected treatment effect for those with particular covariate values. Under certain identifiability assumptions (Dahabreh and Hernán, 2019), the CATE can be estimated parametrically, e.g. via linear regression, or non-parametrically, e.g. via nearest-neighbour matching , kernel methods (Brantner et al., 2023), or tree-based methods such as causal random forests (Wager and Athey, 2018) and their Bayesian counterpart (Hahn et al., 2020).
The approach we follow in this paper to fuse data from an RCT with data from an observational study builds on several lines of work. Dahabreh and Hernán (2019) explored the generalisability of RCT inference to different target populations using g-methods (including reweighting methods) and doubly robust estimators. Others have focused on using RCT data to correct for the possible (hidden) confounding effect present in observational data. For example, Wu and Yang (2022) developed an R-learner and proved consistency and asymptotic efficiency of the estimator under the assumption of overlap between the RCT and observational study. Kallus et al. (2018) employed a two-step approach to correct for hidden confounding in the absence of covariate overlap. Their approach relies on the strong assumption of linearity of the hidden confounding effect term (see Section 3.1) and hinges on achieving parametric identification of this correction term, ultimately yielding a consistent estimate of the CATE over the confounded sample. In a related approach, Yang et al. (2022) relaxed the linearity assumption and established identifiability and efficiency results under more general parametric structural models for the confounding effect. Lin and Evans (2023) take yet another approach and use a power likelihood to reduce the effect of hidden confounding.
@@
Our method, Causal-ICM, employs multi-task GPs (see Alvarez et al. (2012) for a review) to combine information from distinct studies for the purpose of CATE estimation. This builds on work that leverages multi-task GPs for solely observational data, whereby each potential outcome is treated as a distinct task (Alaa and van der Schaar, 2017). GP-SLC is a similar approach that is able to handle hierarchical hidden confounders (Witty et al., 2020). Others have built on this framework to develop a general class of counterfactual multi-task deep kernels models that efficiently estimate causal effects and learn policies by stacking coregionalized Gaussian Processes and Deep Kernels (Caron et al., 2022). GPs have also been employed as a matching tool for causal inference (Huang et al., 2023).
@@ 3 Methodology
We first begin with a brief introduction on data fusion for heterogeneous treatment effect estimation. We then introduce a multi-task GPs framework specifically for this application. Throughout the paper, we use bold-face to denote vectors, e.g. $\mathbf{x}$, and capitals to denote matrices, e.g. $X$.
@@ 3.1 Data Fusion for Heterogenous Treatment Effect Esimation
Data fusion refers to the integration of two or more distinct sources of data. In this work, we focus on combining information from an observational dataset and a RCT (experimental) dataset. We assume that the observational study is representative of the target population. In contrast, we generally assume that unequal participation in the RCT results in an unrepresentative sample. In particular, the distribution of covariates within the RCT may be mean-shifted and more concentrated relative to the observational study. Moreover, we will generally assume that the number of individuals in the observational study is larger than that of the RCT, reflecting real-world differences in sample sizes.
We begin with some notation. Let $S\in\{\text{o},\text{e}\}$ be a study indicator, where ‘e’ stands for experimental and ‘o’ stands for observational. Let $A\in\{0,1\}$ denote the binary treatment indicator, where $A=1$ indicates the treatment arm and $A=0$ the control arm. We assume that the same set of baseline covariates are observed in each study, so that for each individual we observe covariates $\mathbf{x}\in\mathcal{X}^{S}\subseteq\mathbb{R}^{p}$ prior to treatment assignment for $S\in\{\text{o},\text{e}\}$, where $\mathcal{X}^{\text{e}}$ has overlap with $\mathcal{X}^{\text{o}}$ but generally $\mathcal{X}^{\text{o}}\not\subseteq\mathcal{X}^{\text{e}}$. Finally, we observe a continuous outcome $Y\in\mathbb{R}$.
We define the counterfactual outcome $Y_{i}^{a}$ as the outcome we would have observed had the $i^{\text{th}}$ individual been assigned treatment $A=a$, where we assume no interference. The causal estimand of interest is the CATE comparing $A=1$ versus $A=0$ in the target population given the baseline covariates ${X}$, defined as
$\begin{split}\tau(\mathbf{x})=\mathbb{E}\left[Y^{a=1}-Y^{a=0}|\mathbf{x}\right]\end{split}$
@@
for $\mathbf{x}\in\mathcal{X}^{\text{e}}\cup\mathcal{X}^{\text{o}}$. The difference in conditional average outcomes in each study is denoted by
$\omega^{s}(\mathbf{x})=\mathbb{E}\left[Y|\mathbf{x},A=1,S=s\right]-\mathbb{E}\left[Y|\mathbf{x},A=0,S=s\right].$
Under standard identifiability assumptions for the RCT (Appendix A.1), the CATE can be identified from the RCT data as
$\begin{split}\tau(\mathbf{x})=\omega^{\text{e}}(\mathbf{x})\end{split}$
for $\mathbf{x}\in\mathcal{X}^{\text{e}}$. In practice, however, estimating $\omega^{\text{e}}(\mathbf{x})$ even for $\mathbf{x}\in\mathcal{X}^{\text{e}}$ can be challenging in the tails of the covariate distribution for the RCT, where we have fewer samples. Furthermore, we may be interested in a broader target population, which may have support outside of $\mathcal{X}^{\text{e}}$.
@@
On the other hand, the covariate distribution of the observational study may be more representative of the target population compared to that of the RCT. Without the additional unrealistic assumption of no unmeasured confounding in the observational study, however, it is not possible to identify $\tau(\mathbf{x})$ from the observational data alone. In particular, we have that the hidden confounding effect in the observational data denoted by $\eta(\mathbf{x})=\tau(\mathbf{x})-\omega^{\text{o}}(\mathbf{x})$ is nonzero. As a result, estimates of $\omega^{\text{o}}(\mathbf{x})$ are biased for our causal estimand of interest $\tau(\mathbf{x})$.
In our work, we aim to use both sources of information within a multi-task Gaussian process framework to jointly estimate $(\omega^{\text{o}}(\mathbf{x}),\omega^{\text{e}}(\mathbf{x}))$. By treating the two sources of data as different tasks, we aim to take advantage of both the unbiasedness property of the RCT and the support of the observational study. Moreover, by comparing $(\omega^{\text{o}}(\mathbf{x}),\omega^{\text{e}}(\mathbf{x}))$ within $\mathcal{X}^{e}$, we are able to correct for the confounding effect. Unlike Kallus et al. (2018), we do not assume $\eta(\mathbf{x})$ is linear, and instead rely on the multi-task GP’s ability to extrapolate in a non-linear fashion, whilst quantifying uncertainty through the posterior distribution.

### v2

@@ Abstract
Bridging the gap between internal and external validity is crucial for heterogeneous treatment effect estimation. Randomised controlled trials (RCTs), favoured for their internal validity due to randomisation, often encounter challenges in generalising findings due to strict eligibility criteria. Observational studies, on the other hand, may provide stronger external validity through larger and more representative samples but can suffer from compromised internal validity due to unmeasured confounding. Motivated by these complementary characteristics, we propose a novel Bayesian nonparametric approach, Causal-ICM, leveraging multi-task Gaussian processes to integrate data from both RCTs and observational studies. In particular, we introduce a parameter that controls the degree of borrowing between the datasets and prevents the observational dataset from dominating the estimation. We propose a data-adaptive procedure for choosing the optimal value of the parameter. Causal-ICM outperforms other data fusion methods in point estimation across the covariate support of the observational study and provides principled uncertainty quantification for the estimated treatment effects. We demonstrate the robust performance of Causal-ICM in diverse scenarios through multiple simulation studies and a real-world study.
keywords:
Data Fusion, Bayesian Nonparametrics, Multitask Gaussian Processes, Generalisability, Heterogeneous Treatment Effects
@@ 1 Introduction
Treatment effect estimation is an important task in many applications, including medicine, epidemiology, and the social sciences. The goal is to quantify the expected impact of a particular intervention in a given target population. The effect of an intervention within a population may vary systematically with respect to a particular set of covariates. Understanding such treatment effect heterogeneity is critical to ensure that appropriate decisions are taken for all individuals, not just those similar to the average (Brantner et al., 2023). An appropriate characterisation of the uncertainty of treatment effect estimates is also required to support robust and reliable decision-making. Uncertainty quantification for heterogeneous treatment effects is challenging, however, especially when making inferences about individuals not represented in the study. Classical methods typically rely heavily on extrapolation without adequate variance inflation outside the study support (Degtiar and Rose, 2023).
  The gold standard for treatment effect estimation remains the RCT. The random treatment allocation in the study guards against the effect of confounding As a result, under standard identifiability assumptions, we can obtain an unbiased estimate of the treatment effect. There are, however, drawbacks to the use of RCTs (Frieden, 2017), such as the high financial cost or the limited sample sizes. While they may be powered enough to detect the Average Treatment Effect (ATE), they are typically underpowered for subgroup effects. Furthermore, strict eligibility criteria and selection bias may lead to a sample non-representative of the target population (National Academies of Sciences, Engineering, and Medicine, 2022).
@@
Observational studies provide an alternative source of data for treatment effect estimation. In contrast to RCTs, they generally benefit from larger sample sizes and tend to be more representative of the target population. However, observational studies may suffer from unobserved confounding, i.e. unobserved variables that affect both the outcome of interest and the treatment allocation, and can lead to biased treatment effect estimates. For example, a patient’s underlying health status might influence both the treatment they receive and their eventual outcome, yet remain unmeasured.
@@
  The complementary nature of RCT and observational data indicates the potential benefits from combining the two sources to obtain better heterogeneous treatment effect estimates. Such data fusion approaches have received increased interest in recent years. A range of methods have been developed to integrate information from multiple sources of data, including a limited number targeting causal inference problems (Colnet et al. (2023), Lin et al. (2025b)). These typically rely on strong, untestable assumptions about the structure of the confounding effect and, moreover, are often not able to provide uncertainty quantification around the point estimates of heterogeneous treatment effects.
  In this paper, we introduce a Bayesian nonparametric approach to obtain both point and uncertainty estimates of heterogeneous treatment effects for the target population. Our approach is based on a multi-output Gaussian process (GP), a natural choice in the data fusion context as it enables joint modelling of the RCT and observational outcome regression functions, and uses the Bayesian machinery to share information and quantify uncertainty of the functions in regions of differing covariate support. Our key contributions are as follows:
  • We propose Causal-ICM, a multi-task GP model for heterogeneous treatment effect estimation that accurately captures complex functional relationships in the presence of unobserved confounding both within the support of the RCT covariate distribution and when extrapolating beyond it.
  • Causal-ICM provides principled uncertainty quantification across the full support of the target population. Crucially, we theoretically show that Causal-ICM limits the amount of information learnt from the observational dataset, and safeguards against overconfidence in the presence of bias.
  • Through a comprehensive set of simulation studies and an application to a real-world dataset, we demonstrate that Causal-ICM achieves similar or superior performance both in point estimation and uncertainty quantification relative to a broad set of state-of-the-art causal data fusion methods.
@@ 2 Related Literature
Heterogeneous treatment effect (HTE) estimation has received renewed interest in recent years. The most common strategy is to focus on the conditional average treatment effect (CATE) function, which quantifies the expected treatment effect given particular covariate values. Under suitable identifiability assumptions (Dahabreh and Hernán, 2019), the CATE can be estimated parametrically - e.g. using linear regression - or non-parametrically, e.g. using nearest-neighbour matching, kernel methods (Brantner et al., 2023), or tree-based methods including causal random forests (Wager and Athey, 2018) and their Bayesian counterpart (Hahn et al., 2020).
Data fusion across an RCT and an observational study has tended to focus on ATE estimation (see Lin et al. (2024) and Colnet et al. (2023) for comprehensive reviews). These methods aim to improve efficiency by integrating information from both sources. Demirel et al. (2024) focus instead on data fusion for generalisability, proposing a framework that builds an observational predictor and then treat the difference between the trial outcome function and that observational predictor as a bias function that captures both confounding and transportation bias due to the differences between the trial and target populations. Similar approaches, although outside the immediate scope of our work, have focused on data fusion for the estimation of long term effects, where randomised data are unconfounded but contain only short term effects, while observational data, although confounded, contain information about long term outcomes (Ghassami et al. (2025), Imbens et al. (2024)).
Data fusion methods for CATE estimation, the focus of this paper, follow three broad lines of work. One strategy is to use experimental data to debias observational estimates by learning a correction or bias function. Kallus et al. (2018) introduce a two-step approach that corrects hidden confounding in the absence of covariate overlap, relying on the strong assumption of linear confounding and the ability to identify this correction term parametrically. Yang et al. (2022b) relax this linearity assumption and provide identifiability and efficiency results under more flexible structural models. Hatt et al. (2022) develop a representation-learning strategy that learns shared features and confounded outcome models from observational data, using the RCT to estimate a bias function that debiases these models, with finite-sample bounds that highlight how performance depends on sample size, distribution shift, and bias complexity. Wu and Yang (2022) propose an R-learner that achieves consistency and asymptotic efficiency under covariate overlap between the RCT and the observational study. A second direction consists of “test-then-pool” strategies, in which the trial and observational datasets are combined only if the observational study appears sufficiently unbiased. For instance, Yang et al. (2023) implement such a strategy, although their approach is limited to linear treatment effect models and relies on strong parametric assumptions. A third line of work uses Bayesian dynamic borrowing to regulate how much information is extracted from the observational study. For instance, Lin and Evans (2023) propose a Bayesian dynamic borrowing approach through a power likelihood to reduce the effect of confounding and control the degree of information we borrow from the observational study.
@@
Our method, Causal-ICM, takes a Bayesian dynamic borrowing approach based on multi-task GPs, whereby each potential outcome is treated as a distinct task. Multi-task GPs have been successfully employed for causal inference using solely observational data (Alaa and van der Schaar, 2017) and extended to handle hierarchical hidden confounders (Witty et al., 2020). Others have built on this framework to develop a general class of counterfactual multi-task deep kernel models that efficiently estimate causal effects and learn policies by stacking coregionalized GPs and deep kernels (Caron et al., 2022). GPs have also been employed as a matching tool for causal inference (Huang et al., 2023). However, these methods are limited to purely observational settings; here we extend these ideas by enabling inference in a data-fusion setting. Our approach is most similar to concurrent work employing multi-task GPs for pseudo-outcome regression to obtain error bounds on the observational bias term (Fawkes et al., 2025).
@@ 3 Methodology
We begin with a brief introduction on data fusion for HTE estimation. We then introduce a multi-task GPs framework tailored for this purpose. Throughout the paper, we use bold-face to denote vectors, e.g. $\mathbf{x}$, and capitals to denote matrices, e.g. $X$.
@@ 3.1 Data Fusion for Heterogenous Treatment Effect Esimation
In this work, we focus on combining information from an observational dataset and a RCT (experimental) dataset. We begin with some notation. Let $S\in\{\text{o},\text{e}\}$ indicate the study, with ‘e’ denoting the experimental sample and ‘o’ the observational one. Let $A\in\{0,1\}$ be the binary treatment indicator. We observe the same baseline covariates in both studies, so each individual has $\mathbf{X}\in\mathcal{X}^{S}\subseteq\mathbb{R}^{p}$, where $\mathcal{X}^{\text{e}}$ and $\mathcal{X}^{\text{o}}$ overlap but $\mathcal{X}^{\text{o}}$ is not necessarily contained in $\mathcal{X}^{\text{e}}$. Without loss of generality we divide the baseline covariates into three subsets, the treatment effect modifiers $\mathbf{X}_{\tau}\in\mathcal{X}_{\tau}^{\text{e}}$, the confounders, $\mathbf{X}_{W}\in\mathcal{X}_{W}^{\text{e}}$, and the selection variables, $\mathbf{X}_{S}\in\mathcal{X}_{S}^{\text{e}}$, such that $\mathbf{X}=\mathbf{X}_{\tau}\cup\mathbf{X}_{W}\cup\mathbf{X}_{S}$ Finally, we observe a continuous outcome $Y\in\mathbb{R}$. The counterfactual outcome $Y_{i}^{a}$ denotes the value individual $i$ would have had under treatment $A=a$. The causal estimand of interest is the CATE comparing $A=1$ versus $A=0$ in the target population given the baseline covariates X{X}, defined as
$\begin{split}\tau(\mathbf{x}_{\tau})=\mathbb{E}\left[Y^{a=1}-Y^{a=0}|\mathbf{X}_{\tau}=\mathbf{x}_{\tau}\right]\end{split}$
@@
for $\mathbf{x}_{\tau}\in\mathcal{X}^{\text{e}}\cup\mathcal{X}^{\text{o}}$. The difference in the conditional expectations of the outcomes in each study is
$\omega^{s}(\mathbf{x}_{\tau})=\mathbb{E}\left[Y|\mathbf{X}_{\tau}=\mathbf{x}_{\tau},A=1,S=s\right]-\mathbb{E}\left[Y|\mathbf{X}_{\tau}=\mathbf{x}_{\tau},A=0,S=s\right].$
The fundamental problem of causal inference is that for each unit we cannot observe both potential outcomes simultaneously. Hence, we make the following identifiability assumptions, which are standard in the literature (e.g. Lanners et al. (2025), Shi et al. (2023)):
  1. A1
    Consistency: $A=a\Rightarrow Y=Y^{a}$
  2. A2
    Mean conditional exchangeability over treatment for the treatment contrast given $\mathbf{X}_{\tau}$ in the RCT
    $\mathbb{E}[Y^{1}-Y^{0}\mid\mathbf{X}_{\tau},S=\text{e}]$ $=\mathbb{E}[Y^{1}\mid A=1,\mathbf{X}_{\tau},S=\text{e}]-\mathbb{E}[Y^{0}\mid A=0,\mathbf{X}_{\tau},S=\text{e}]$
  3. A3
    Positivity of treatment assignment in the RCT: $P(A=a|\mathbf{X}_{S}=\mathbf{x}_{S},S=\text{e})>0$ for $\mathbf{x}_{S}\in\mathcal{X}_{S}^{\text{e}}$.
  4. A4
    Mean conditional exchangeability over selection into the RCT for the treatment contrast given $\mathbf{X}_{\tau}$, $\mathbf{X}_{S}$:
    $\mathbb{E}[Y^{1}-Y^{0}\mid\mathbf{X}_{\tau}\cup\mathbf{X}_{S}]=\mathbb{E}[Y^{1}-Y^{0}\mid\mathbf{X}_{\tau}\cup\mathbf{X}_{S},S=\text{e}]$
  5. A5
    Positivity of RCT participation: $P(S=\text{e}|\mathbf{X}_{S}=\mathbf{x}_{S})>0$ for $\mathbf{x}_{S}\in\mathcal{X}^{\text{e}}_{S}$
Under assumptions A1 - A5 the CATE can be identified from the RCT data as
$\tau(\mathbf{x}_{\tau})=\omega^{\text{e}}(\mathbf{x}_{\tau})$
for $\mathbf{x}_{\tau}\in\mathcal{X}_{\tau}^{\text{e}}$. The derivation is provided in the Appendix A. For notational simplicity, and with a slight abuse of notation, we henceforth use $\mathbf{X}$ and $\mathbf{x}$ to denote any subset of covariates, whether treatment-effect modifiers or selection covariates.
@@
Conversely, relative to the RCT, the covariate distribution of the observational study may be more representative of the target population. In the presence of unmeasured confounding, however, it is not possible to identify $\tau(\mathbf{x})$ from the observational data alone. In particular, the hidden confounding effect in the observational data, denoted by $\eta(\mathbf{x})=\tau(\mathbf{x})-\omega^{\text{o}}(\mathbf{x})$, is nonzero. As a result, estimates of $\omega^{\text{o}}(\mathbf{x})$ may be biased for the estimand $\tau(\mathbf{x})$.
In this work, we use a multi-task Gaussian process to jointly estimate $(\omega^{\text{o}}(\mathbf{x}),\omega^{\text{e}}(\mathbf{x}))$, treating the two data sources as separate tasks. This leverages both the unbiasedness of the RCT and the broader covariate support of the observational study, while adaptively accounting for confounding by comparing $(\omega^{\text{o}}(\mathbf{x}),\omega^{\text{e}}(\mathbf{x}))$ within $\mathcal{X}^{\text{e}}$. Unlike other methods, we avoid strong parametric assumptions about the bias function, relying instead on the multi-task GP to extrapolate non-linearly with posterior uncertainty quantification, so that regions with low covariate support naturally exhibit higher uncertainty.